OpenStudy (tomfoolery1):

Suppose that four students scores are selected randomly from the exam that has a st dev of 1.8 and a mean of 70 what is the probability that the average score for those four selected exams is greater than 73?

jimthompson5910 (jim_thompson5910):

You're working with the xbar distribution (distribution of sample means) \(\Large \bar{x}\) = xbar = 73 \(\Large \mu\)= mu = 70 \(\Large \sigma\)= sigma = 1.8 n = 4 Test Statistic \[\Large z = \frac{\bar{x} - \mu}{\frac{\sigma}{\sqrt{n}}}\] \[\Large z = \frac{73-70}{\frac{1.8}{\sqrt{4}}}\] \[\Large z = ??\] Tell me what you get

jimthompson5910 (jim_thompson5910):

let me know if you get stuck @Tomfoolery1

OpenStudy (tomfoolery1):

3.3333? i hope thats right

jimthompson5910 (jim_thompson5910):

yep that's the correct z score now use either a calculator or a table to compute P(Z > 3.33)

jimthompson5910 (jim_thompson5910):

The table should be in the back of your text book

OpenStudy (tomfoolery1):

0.999566 this is what i got

jimthompson5910 (jim_thompson5910):

http://www.stat.ufl.edu/~athienit/Tables/Ztable.pdf using that table P(Z < 3.33) = 0.9996 so P(Z > 3.33) = 1 - P(Z < 3.33) P(Z > 3.33) = 1 - 0.9996 P(Z > 3.33) = 0.0004

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):